# 5- V o l u m e 1, N u m b e r 1, S e p t e m b e lr 2 0 1 8

### v2

@@ 3.1 DoF analysis
$\mathbf{y}_{k}=(\mathbf{\underbrace{\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}}_{\mathbf{A}}}+\mathbf{H}_{dk})\mathbf{x}+\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{n}+\mathbf{z}.$ (9)
@@
$\mathbf{\hat{H}}$ $=\mathbf{\underbrace{\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}}_{\mathbf{A}}}+\mathbf{H}_{dk}.$ (10)
@@
$\leq$ $\mathrm{rank}\left(\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\right)+\mathrm{rank}\left(\mathbf{H}_{dk}\right).$ (12)
@@
$\mathrm{rank}\left(\mathbf{AB}\right)\leq\mathrm{min}\left(\mathrm{rank}\left(\mathbf{A}\right),\left(\mathrm{rank}\mathbf{B}\right)\right),$ (13)
@@
$\mathrm{rank}\left(\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\right)$
$\leq\mathrm{min}\left(\mathrm{rank}\left(\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\right),\mathrm{rank}\left(\mathbf{H}_{BI}\right)\right)$
$\leq\mathrm{min}\left(\mathrm{rank}\left(\mathbf{G}_{k}^{H}\right),\mathrm{rank}\left(\mathbf{\Theta}_{k}\right),\mathrm{rank}\left(\mathbf{H}_{BI}\right)\right).$ (14)
Considering the matrix $\mathbf{\Theta}_{k}$ is a full-rank diangonal matrix, the above inequality reduces to
$\mathrm{rank}\left(\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\right)\leq\mathrm{min}(\mathrm{rank}(\mathbf{G}_{k}^{H}),\mathrm{rank}(\mathbf{H}_{BI})).$ (15)
@@
$\mathrm{rank}\left(\mathbf{H}_{dk}+\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\right)\leq 2,$ (18)
@@
In rank-deficient channel,
@@
where all elements of matrices $\mathbf{G}_{k}^{H}$, $\mathbf{H}_{dk}$ and $\mathbf{H}_{BI}$ are independent and identically distributed (iid), so we have the following inequality
$\mathrm{rank}\left(\mathbf{H}_{dk}+\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\right)\leq\mathrm{min}(Q_{k},M)$
$+\mathrm{min}\left(\mathrm{min}(N,Q_{k}),\mathrm{min}(M,N)\right).$ (20)
@@
$\mathbf{y}_{k}=$ $\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\right)\mathbf{H}_{BI}\right)\mathbf{x}$
$+$ $\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{n}+\mathbf{z}.$ (21)
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\right)$
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\right)\mathbf{H}_{BI}\right)$ (24)
@@
$\mathrm{rank}\left(\mathbf{G}_{1}^{H}\right)=\dots=\mathrm{rank}\left(\mathbf{G}_{k}^{H}\right)=\dots=\mathrm{rank}\left(\mathbf{G}_{k}^{H}\right)$
$=1,$ (25)
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\right)\mathbf{H}_{BI}\right)\leq K+1.$ (26)
@@
where $\mathbf{G}_{k}^{H}$ and $\mathbf{H}_{dk}$ are rank-deficient and independent with each other with $Q_{k}\geq I_{k}+J_{k}$ due to the fact that there are a few reflecting or scattering clusters, i.e,, no rich scattering environment, so they have the following inequality
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\right)$
@@
which gives the achievable upper bound of IRS-aided MU-MIMO systems. In other words, the bound will determine the maximum number of bit streams parallelly transmitted from BS to all users with the aid of IRS.
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\right)\mathbf{H}_{BI}\right)\leq\sum_{k=1}^{K}I_{k}+\sum_{k=1}^{K}J_{k},$ (30)
with equality if and only if $\mathrm{range}(\mathbf{H}_{dk})\cap\mathrm{range}(\mathbf{G}_{k}^{H})=\{0\}$ and $\mathrm{range}(\mathbf{H}_{dk}^{H})\cap\mathrm{range}(\mathbf{H}_{BI}^{H}\mathbf{\Theta}_{k}^{H}\mathbf{G}_{k})=\{0\}$.
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\right)\leq 2\sum_{k=1}^{K}I_{k}.$ (31)
@@
Observing the above two results, it very clear that compared with the no-IRS MU-MIMO network, the DoF of the IRS-aided MU-MIMO network doubles in a severe rank-deficient channel due to the help of IRS.
Now, we consider an extreme case of low-rank, when all $\mathrm{rank}(\mathbf{H}_{dk})=0$, this means that all channels from BS to users are blocked. If $\mathrm{rank}(\mathbf{G}^{H}_{k})=J$ and $\mathrm{rank}(\mathbf{H}_{BI})=\mathrm{min}\{M,N\}$, where all $\mathbf{G}^{H}_{k}$ are iid, then the total DoF of system grows from 0 to $\mathrm{min}\{KJ,M,N\}$, where J is the number of antennas per user. Such a DoF enhancement is very dramatic as $K$, $M$ and $N$ goes to large-scale. Similarly, when all $\mathrm{rank}(\mathbf{H}_{dk})=1$, this means all $\mathbf{H}_{dk}$ are LoS channels, we will attain the same result. This means IRS will play an extremely important role in DoF increasement in a low-rank channel. In other words, a low-rank scenario needs more IRS compared to the rich-scattered environment. Here, a rich-scattered means all channel matrices $\mathrm{rank}(\mathbf{G}^{H}_{k})$, $\mathrm{rank}(\mathbf{H}_{dk})$ and $\mathrm{rank}(\mathbf{H}_{BI})$ are full-rank and iid. We may make a conclusion that there is no space to improve DoF even with the help of IRS in such a full-rank channel in accordance with inequality (24).
@@ 3.2 Condition for achieving the maximum DoF in LoS channel
$=\mathbf{h}^{H}(\theta_{Bt})\mathbf{h}(\theta_{Bkt})\bullet\mathbf{h}(\theta_{Br})\mathbf{h}^{H}(\theta_{Bkr}).$ (37)
@@
$\cos(\theta_{Bt})+\cos(\theta_{Bkt})$ $=\frac{\lambda i}{Md},$ (39)
@@
$\theta_{Bt}=\arccos(-\cos(\theta_{Bkt})+\frac{\lambda i}{Md}).$ (40)
@@
$\cos(\theta_{Bt})=\frac{\lambda i}{Md}-\frac{1}{2}.$ (41)
@@ 4 Proposed Beamforming Methods
In order to reduce the interference form multi-stream and multi-user transmission, in this section, the three low-complexity closed-form beamformers, called TLL-MMSE, NSP-MTP-MRP and SO-MMSE are proposed.
@@ 4.1 Proposed TLL-MMSE in rank-deficient channel
$+\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\sum_{i=1,i\neq k}^{K}\mathbf{V}_{i}\mathbf{s}_{i}+\mathbf{U}_{k}^{H}\sum_{j=1}^{K}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{n},$ (43)
where $\mathbf{G}^{H}_{jk}\in\mathbb{C}^{Q_{k}\times N_{k}}$ denotes the channel gain matrix from IRS $j$ to user $k$.
Then the rate of user $k$ can be obtained
$\mathbf{\gamma}_{k}$ $=\log_{2}\left|\mathbf{I}_{k}+\frac{\mathbf{A}_{1}+\mathbf{A}_{2}}{\mathbf{B}_{1}+\mathbf{B}_{2}+\mathbf{B}_{3}}\right|,$ (44)
@@
$\mathbf{B}_{3}$ $=\sum_{j=1}^{K}\sigma^{2}\mathbf{U}_{k}^{H}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{\Theta}_{j}^{H}\mathbf{G}_{jk}\mathbf{U}_{k}+\sigma^{2}_{z}\mathbf{U}_{k}^{H}\mathbf{U}_{k}.$ (45)
@@
where $P_{Tk}$ represents the transmit power that BS allocates to $\mathbf{V}_{k}$, and
@@
$B_{3}$ $=\sigma^{2}+\sigma^{2}_{z}.$ (48)
Let us define $\mathbf{A}=\mathbf{H}_{dk}^{H}\mathbf{H}_{dk}+\mathbf{H}_{BIk}^{H}\mathbf{H}_{BIk}$, $\mathbf{B}=\sum_{j=1,j\neq k}^{K}(\mathbf{H}_{dj}^{H}\mathbf{H}_{dj}+\mathbf{H}_{BIj}^{H}\mathbf{H}_{BIj})$ and $C=\sigma^{2}+\sigma^{2}_{z}$. Defining $\mathbf{V}_{k}=\sqrt{P_{Tk}}\hat{\mathbf{V}}_{k}$, then $SLNR_{k}$ can be reduced to the following form
@@
$\max_{\mathbf{\Theta}_{k}}SLNR_{k}=\frac{\|\mathbf{G}_{kk}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BIk}\mathbf{V}_{k}\mathbf{s}_{k}\|^{2}}{\sum_{j=1,j\neq k}^{K}\|\mathbf{G}_{kj}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BIk}\mathbf{V}_{k}\mathbf{s}_{k}\|^{2}+\sigma^{2}_{z}}$
@@
Let us define $\mathbf{\Theta}_{k}=\text{diag}(\boldsymbol{\theta}_{k})$, hence (4.1) can be rewritten as
$\max_{\mathbf{\Theta}_{k}}SLNR_{k}=\frac{\boldsymbol{\theta}_{k}^{H}\mathbf{A}^{H}\mathbf{G}_{kk}\mathbf{G}_{kk}^{H}\mathbf{A}\boldsymbol{\theta}_{k}}{\sum_{j=1,j\neq k}^{K}\boldsymbol{\theta}_{k}^{H}\mathbf{A}^{H}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H}\mathbf{A}\boldsymbol{\theta}_{k}+\sigma^{2}_{z}}$
@@
where $\mathbf{A}=\text{diag}(\mathbf{H}_{BIk}\mathbf{V}_{k}\mathbf{s}_{k})$. Assuming $P_{Ik}=\|\mathbf{\Theta}_{k}\mathbf{H}_{BIk}\mathbf{V}_{k}\mathbf{s}_{k}\|^{2}+\|\mathbf{\Theta}_{k}\|^{2}_{F}\sigma^{2}$, and $\boldsymbol{\theta}_{k}=\hat{\boldsymbol{\theta}}_{k}\hat{\mu}_{k}$, where $\hat{\mu}_{k}=\|\boldsymbol{\theta}_{k}\|$, $\hat{\boldsymbol{\theta}}_{k}^{H}\hat{\boldsymbol{\theta}}_{k}=1$. The (4.1) can be rewritten as
$\max_{\hat{\boldsymbol{\theta}}_{k}}SLNR_{k}$
$=$ $\frac{\hat{\boldsymbol{\theta}}_{k}^{H}\mathbf{A}^{H}\mathbf{G}_{kk}\mathbf{G}_{kk}^{H}\mathbf{A}\hat{\boldsymbol{\theta}}_{k}}{\hat{\boldsymbol{\theta}}_{k}^{H}(\sum_{j=1,j\neq k}^{K}\mathbf{A}^{H}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H}\mathbf{A}+\frac{\sigma^{2}_{z}}{\hat{\mu}_{k}^{2}}\mathbf{I}_{k})\hat{\boldsymbol{\theta}}_{k}}.$ (53)
Then (4.1) can be approximated by the following
$\max_{\hat{\boldsymbol{\theta}}_{k}}SLNR_{k}$
$=$ $\frac{\hat{\boldsymbol{\theta}}_{k}^{H}\mathbf{A}^{H}\mathbf{G}_{kk}\mathbf{G}_{kk}^{H}\mathbf{A}\hat{\boldsymbol{\theta}}_{k}}{\hat{\boldsymbol{\theta}}_{k}^{H}(\sum_{j=1,j\neq k}^{K}\mathbf{A}^{H}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H}\mathbf{A}+\sigma^{2}_{z}\mathbf{I}_{k})\hat{\boldsymbol{\theta}}_{k}},$ (54)
@@
$\hat{\mathbf{D}}_{k}=\frac{\mathbf{A}^{H}\mathbf{G}_{kk}\mathbf{G}_{kk}^{H}\mathbf{A}}{\sum_{j=1,j\neq k}^{K}\mathbf{A}^{H}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H}\mathbf{A}+\sigma^{2}_{z}\mathbf{I}_{k}}.$ (55)
@@
$\hat{\mu}$ $=\sqrt{\frac{P_{Ik}}{\hat{\boldsymbol{\theta}}_{k}^{H}(\mathbf{A}^{H}\mathbf{A}+\sigma^{2}\mathbf{I}_{k})\hat{\boldsymbol{\theta}}_{k}}},$ (56)
@@
$+\sum_{j=1}^{K}\sigma^{2}\mathbf{U}_{k}^{H}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{\Theta}_{j}^{H}\mathbf{G}_{jk}\mathbf{U}_{k}+\sigma^{2}_{z}\mathbf{U}_{k}^{H}\mathbf{U}_{k}.$ (58)

### v3

@@ 3.1 DoF analysis
$\mathbf{y}_{k}=(\mathbf{\underbrace{\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}}_{\mathbf{A}}}+\mathbf{H}_{dk})\mathbf{x}+\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{n}+{\mathbf{z}_{k}}.$ (9)
@@
$\mathbf{\hat{H}}$ $=\mathbf{\underbrace{\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}}_{\mathbf{A}}}+\mathbf{H}_{dk}.$ (10)
@@
$\leq$ $\mathrm{rank}\left(\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\right)+\mathrm{rank}\left(\mathbf{H}_{dk}\right).$ (12)
@@
$\mathrm{rank}\left(\mathbf{AB}\right)\leq\mathrm{min}\left(\mathrm{rank}\left(\mathbf{A}\right),\mathrm{rank}\left(\mathbf{B}\right)\right),$ (13)
@@
$\mathrm{rank}\left(\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\right)$
$\leq\mathrm{min}\left(\mathrm{rank}(\mathbf{G}_{k}^{H}\mathbf{\Theta}),\mathrm{rank}\left(\mathbf{H}_{BI}\right)\right)$
$\leq\mathrm{min}\left(\mathrm{rank}(\mathbf{G}_{k}^{H}),\mathrm{rank}\left(\mathbf{\Theta}\right),\mathrm{rank}\left(\mathbf{H}_{BI}\right)\right).$ (14)
Considering the matrix $\mathbf{\Theta}$ is a full-rank diangonal matrix, the above inequality reduces to
$\mathrm{rank}\left(\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\right)\leq\mathrm{min}(\mathrm{rank}(\mathbf{G}_{k}^{H}),\mathrm{rank}(\mathbf{H}_{BI})).$ (15)
@@
$\mathrm{rank}\left(\mathbf{H}_{dk}+\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\right)$
@@
$\mathrm{rank}\left(\mathbf{H}_{dk}+\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\right)\leq 2,$ (18)
@@
In full-rank channel,
@@
where all elements of matrices $\mathbf{G}_{k}^{H}$, $\mathbf{H}_{dk}$ and $\mathbf{H}_{BI}$ are independent and identically distributed (i.i.d.), so we have the following inequality
$\mathrm{rank}\left(\mathbf{H}_{dk}+\mathbf{G}_{k}^{H}\mathbf{{\color[rgb]{0,0,0}\Theta}}\mathbf{H}_{BI}\right)\leq\mathrm{min}(M,Q_{k})$
$+\mathrm{min}\left(\mathrm{min}(N,Q_{k}),\mathrm{min}(N,M)\right).$ (20)
@@
$\mathbf{y}_{k}=$ $\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{{\color[rgb]{0,0,0}\Theta}}\right)\mathbf{H}_{BI}\right)\mathbf{x}$
$+$ $\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{{\color[rgb]{0,0,0}\Theta}}\mathbf{n}+\mathbf{z}.$ (21)
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{{\color[rgb]{0,0,0}\Theta}}\mathbf{H}_{BI}\right)$
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{{\color[rgb]{0,0,0}\Theta}}\right)\mathbf{H}_{BI}\right)$ (24)
@@
$\mathrm{rank}\left(\mathbf{H}_{d1}\right)=\dots=\mathrm{rank}\left(\mathbf{H}_{dK}\right)=1,$ (25)
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{dk}+\left(\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{{\color[rgb]{0,0,0}\Theta}}\right)\mathbf{H}_{BI}\right)\leq K+1.$ (26)
@@
where $\mathbf{G}_{k}^{H}$ and $\mathbf{H}_{dk}$ are rank-deficient and independent with each other with $Q_{k}\geq I_{k}+J_{k}$, due to the fact that there are a few reflecting or scattering clusters, i.e., no rich-scattered environment, so they have the following inequality
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{k}\right)$
@@
which gives the achievable upper bound of IRS-aided MU-MIMO systems, where $\mathbf{H}_{k}=\mathbf{H}_{dk}+\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}$. In other words, the bound will determine the maximum number of bit streams parallelly transmitted from BS to all users with the aid of IRS.
Considering that $\mathbf{H}_{k}$ is a $Q_{k}\times M$ channel matrix, so
$\mathrm{rank}\left(\mathbf{H}_{k}\right)\leq\mathrm{min}\left(Q_{k},M\right),$ (30)
where $M\gg Q_{k}$, i.e. $\mathrm{rank}\left(\mathbf{H}_{k}\right)\leq\mathrm{min}\left(Q_{k}\right)$.
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}{\color[rgb]{0,0,0}\mathbf{H}_{k}}\right)\leq\sum_{k=1}^{K}I_{k}+\sum_{k=1}^{K}J_{k},$ (31)
with equality if and only if $\mathrm{range}(\mathbf{H}_{dk})\cap\mathrm{range}(\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI})=\{0\}$ and $\mathrm{range}(\mathbf{H}_{dk}^{H})\cap\mathrm{range}(\mathbf{H}_{BI}^{H}\mathbf{\Theta}^{H}\mathbf{G}_{k})=\{0\}$.
@@
$\mathrm{rank}\left(\sum_{k=1}^{K}{\color[rgb]{0,0,0}\mathbf{H}_{k}}\right)\leq 2\sum_{k=1}^{K}I_{k}.$ (32)
@@
Observing the above two results, it very clear that compared with no-IRS MU-MIMO network, the upper bound of DoF of the IRS-aided MU-MIMO network doubles in a rank-deficient channel due to the help of IRS.
Specially, $I_{k}=1$ and $J_{k}=Q_{k}$, i.e., $\mathbf{H}_{dk}$ and $\sum_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}$ are low-rank and full-rank channels, respectively. Then (31) can be written as
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{k}\right)\leq\sum_{k=1}^{K}Q_{k}.$ (34)
Without the aid of IRS, we have
$\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{H}_{dk}\right)\leq K,$ (35)
it is known that the upper bound of DoF of the IRS-aided MU-MIMO network is $(\sum_{k=1}^{K}Q_{k})/K$ times that of no IRS a rank-deficient channel due to the help of IRS.
@@ 3.2 Condition for achieving the maximum DoF in LoS channel
$=\left(\mathbf{h}^{H}(\theta_{Bt})\mathbf{h}(\theta_{Bkt})\right)\left(\mathbf{h}(\theta_{Br})\mathbf{h}^{H}(\theta_{Bkr})\right).$ (40)
@@
$\cos(\theta_{Bt})-\cos(\theta_{Bkt})$ $=\frac{\lambda i}{Md},$ (42)
@@
$\theta_{Bt}=\arccos(\cos(\theta_{Bkt})+\frac{\lambda i}{Md}).$ (43)
@@
$\cos(\theta_{Bt})=\frac{\lambda i}{Md}+\frac{1}{2}.$ (44)
@@ 4 Proposed Beamforming Methods
In order to reduce the interference from multi-stream and multi-user transmission, in this section, the three low-complexity closed-form beamformers, called TLL-MMSE, NSP-MTP-MRP and SO-MMSE are proposed.
@@ 4.1 Proposed TLL-MMSE in rank-deficient channel
$+\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\sum_{i=1,i\neq k}^{K}\mathbf{V}_{i}\mathbf{s}_{i}+\mathbf{U}_{k}^{H}\sum_{j=1}^{K}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{n}_{k},$ (46)
where $\mathbf{G}^{H}_{jk}\in\mathbb{C}^{Q_{k}\times N_{k}}$ denotes the channel gain matrix from IRS $j$ to user $k$, $\mathbf{n}_{k}$ denotes the additive white Gaussian noise (AWGN) at $k$-th IRS , and $\mathbf{n}_{k}\sim\mathcal{CN}(0,{\color[rgb]{0,0,0}\sigma^{2}_{irs,k}}\mathbf{I}_{N})$.
Then the corresponding rate can be given by
$\mathbf{\gamma}_{k}$ $=\log_{2}\left|\mathbf{I}_{k}+(\mathbf{A}_{1}+\mathbf{A}_{2})(\mathbf{B}_{1}+\mathbf{B}_{2}+\mathbf{B}_{3})^{-1}\right|,$ (47)
@@
$\mathbf{B}_{3}$ $=\sum_{j=1}^{K}{\color[rgb]{0,0,0}\sigma^{2}_{irs,k}}\mathbf{U}_{k}^{H}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{\Theta}_{j}^{H}\mathbf{G}_{jk}\mathbf{U}_{k}+\sigma^{2}_{z}\mathbf{U}_{k}^{H}\mathbf{U}_{k}.$ (48)
@@
where $P_{Tk}$ represents the transmit power that BS allocates to $\mathbf{V}_{k}$, and $P_{T}$ is the total transmit power at BS, $P_{T}=\sum_{k=1}^{K}P_{Tk}$, then
@@
$B_{3}$ $={\color[rgb]{0,0,0}\sigma_{irs,k}^{2}}+\sigma^{2}_{z}.$ (51)
Let us define $\mathbf{A}=\mathbf{H}_{dk}^{H}\mathbf{H}_{dk}+\mathbf{H}_{BIk}^{H}\mathbf{H}_{BIk}$, $\mathbf{B}=\sum_{j=1,j\neq k}^{K}(\mathbf{H}_{dj}^{H}\mathbf{H}_{dj}+\mathbf{H}_{BIj}^{H}\mathbf{H}_{BIj})$ and $C={\color[rgb]{0,0,0}\sigma^{2}_{irs,k}}+\sigma^{2}_{z}$. Defining $\mathbf{V}_{k}=\sqrt{P_{Tk}}\hat{\mathbf{V}}_{k}$, then $SLNR_{k}$ can be reduced to the following form
@@
$\max_{\mathbf{\Theta}_{k}}SLNR_{k}=$
$\frac{\|\mathbf{G}_{kk}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BIk}\mathbf{V}_{k}\mathbf{s}_{k}\|^{2}}{\sum_{j=1,j\neq k}^{K}\|\mathbf{G}_{kj}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BIk}\mathbf{V}_{k}\mathbf{s}_{k}\|^{2}+\sum_{j=1}^{K}\|\mathbf{G}_{kj}^{H}\mathbf{\Theta}_{k}\|^{2}\sigma^{2}_{irs,k}}$
@@
where $P_{Ik}$ represents the power at the $k$-th small IRS, and $P_{I}$ is the total power at IRS, $P_{I}=\sum_{k=1}^{K}P_{Ik}$. Let us define $\mathbf{\Theta}_{k}=\text{diag}(\boldsymbol{\theta}_{k})$, hence (4.1) can be rewritten as
$\max_{\mathbf{\Theta}_{k}}SLNR_{k}=$
$\frac{\boldsymbol{\theta}_{k}^{H}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}^{H}\mathbf{G}_{kk}\mathbf{G}_{kk}^{H}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}\boldsymbol{\theta}_{k}}{\boldsymbol{\theta}_{k}^{H}(\sum_{j=1,j\neq k}^{K}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}^{H}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}+\sum_{j=1}^{K}{\color[rgb]{0,0,0}\sigma^{2}_{irs,k}}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H})\boldsymbol{\theta}_{k}}$
@@
where ${\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}=\text{diag}(\mathbf{H}_{BIk}\mathbf{V}_{k}\mathbf{s}_{k})$. Assuming $P_{Ik}=\|\mathbf{\Theta}_{k}\mathbf{H}_{BIk}\mathbf{V}_{k}\mathbf{s}_{k}\|^{2}+\|\mathbf{\Theta}_{k}\|^{2}_{F}{\color[rgb]{0,0,0}\sigma_{irs,k}^{2}}$, and $\boldsymbol{\theta}_{k}=\hat{\boldsymbol{\theta}}_{k}\hat{\mu}_{k}$, where $\hat{\mu}_{k}=\|\boldsymbol{\theta}_{k}\|$, $\hat{\boldsymbol{\theta}}_{k}^{H}\hat{\boldsymbol{\theta}}_{k}=1$. The (4.1) can be rewritten as
$\max_{\hat{\boldsymbol{\theta}}_{k}}SLNR_{k}=$
$\frac{\hat{\boldsymbol{\theta}}_{k}^{H}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}^{H}\mathbf{G}_{kk}\mathbf{G}_{kk}^{H}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}\hat{\boldsymbol{\theta}}_{k}}{\hat{\boldsymbol{\theta}}_{k}^{H}(\sum_{j=1,j\neq k}^{K}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}^{H}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}+\sum_{j=1}^{K}\sigma^{2}_{irs,k}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H})\hat{\boldsymbol{\theta}}_{k}},$ (56)
@@
$\hat{\mathbf{D}}_{k}=\frac{{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}^{H}\mathbf{G}_{kk}\mathbf{G}_{kk}^{H}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}}{\sum_{j=1,j\neq k}^{K}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}^{H}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}+\sum_{j=1}^{K}\sigma^{2}_{irs,k}\mathbf{G}_{kj}\mathbf{G}_{kj}^{H}}.$ (57)
@@
$\hat{\mu}$ $=\sqrt{\frac{P_{Ik}}{\hat{\boldsymbol{\theta}}_{k}^{H}({\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}^{H}{\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathbf{A}_{\theta}}}+{\color[rgb]{0,0,0}\sigma^{2}_{irs,k}}\mathbf{I}_{k})\hat{\boldsymbol{\theta}}_{k}}},$ (58)
@@
$+\sum_{j=1}^{K}{\color[rgb]{0,0,0}\sigma^{2}_{irs,k}}\mathbf{U}_{k}^{H}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{\Theta}_{j}^{H}\mathbf{G}_{jk}\mathbf{U}_{k}+\sigma^{2}_{z}\mathbf{U}_{k}^{H}\mathbf{U}_{k},$ (60)